# Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning

### v2

@@
Abstract.
The growing deployment of large Vision-Language Models (VLMs) exposes critical safety gaps in their alignment mechanisms. While existing jailbreak studies primarily focus on VLMs’ susceptibility to harmful instructions, we reveal a fundamental yet overlooked vulnerability: toxic text continuation – where VLMs produce highly toxic completions when prompted with harmful text prefixes paired with semantically adversarial images. To systematically study this threat, we propose Red Team Diffuser (RTD), the first red teaming diffusion model that coordinates adversarial image generation and toxic continuation through reinforcement learning. Our key innovations include dynamic cross-modal attack and stealth-aware optimization. For toxic text prefixes from an LLM safety benchmark, we conduct greedy search to identify optimal image prompts that maximally induce toxic completions. The discovered image prompts then drive RL-based diffusion model fine-tuning, producing semantically-aligned adversarial images that boost toxicity rates. Stealth-aware optimization introduces joint adversarial rewards that balance toxicity maximization (via Detoxify classifier) and stealthiness (via BERTScore), circumventing traditional noise-based adversarial patterns. Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% over text-only baselines on the original attack set and 8.91% on unseen set, proving generalization capability. Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA. Our findings expose two critical flaws in current VLM alignment: (1) failure to prevent toxic continuation from harmful prefixes, and (2) overlooking cross-modal attack vectors. These results necessitate a paradigm shift toward multimodal red teaming in safety evaluations. We will publicly release the RTD framework to facilitate future research on robust VLM alignment.
Disclaimer: This paper contains content that may be disturbing or offensive.
Red Teaming, Vision-Language Models, Diffusion Models, Reinforcement Learning
ccs: Computing methodologies Artificial intelligence
1. Introduction
Figure 1. Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as general-purpose diffusion model (Rombach et al., 2022).
The rapid deployment of large Vision-Language Models (VLMs), such as GPT-4o (Achiam et al., 2023) and Gemini (Team et al., 2023), has exposed critical blind spots in their multimodal alignment mechanisms. While existing safety strategies predominantly focus on preventing responses to explicit harmful instructions (e.g., ”How to build a bomb?”) through techniques like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), they often fail to address the risks posed by toxic text continuation tasks - where VLMs amplify harmful content when completing partially toxic text prompts paired with semantically adversarial images (Figure 1). This vulnerability originates from two research gaps: (a) Current red teaming methods focus on explicit instruction attacks rather than contextual toxicity induction. (b) Existing safety protocols neglect cross-modal attack vectors where images amplify output toxicity.
The multimodal nature of VLMs introduces unique safety challenges: identical text inputs paired with different images can yield drastically different toxicity levels (Liu et al., 2023), sometimes in unintuitive ways. For instance, Li et al. (Li et al., 2024) demonstrated that pairing jailbreak prompts with even a blank image can significantly increase attack success rates. However, systematically identifying images that effectively trigger toxic continuations remains an open and challenging problem. Recent progress in white-box jailbreak techniques (Qi et al., 2024; Wang et al., 2024b) has explored the use of adversarial images optimized directly against model parameters. While powerful, such gradient-based approaches are impractical in real-world scenarios, particularly within black-box VLM-as-a-Service contexts where full model access is unavailable. Additionally, adversarial images generated through gradient-based methods typically exhibit noticeable distortions, making them susceptible to detection and defense mechanisms (Nie et al., 2022; Zhang et al., 2023; Zhao et al., 2024). Moreover, studies have demonstrated that these adversarial images typically show limited transferability between VLMs, even after successfully jailbreaking multiple VLMs simultaneously (Schaeffer et al., [n. d.]).
These limitations motivate us to develop stealthy black-box attacks that produce natural adversarial images without full model access. Existing black-box jailbreak methods typically adopt suboptimal paradigms, rely on manually engineered pipelines involving typographic manipulations or pre-trained diffusion models to generate query-relevant images (Gong et al., 2023; Liu et al., 2023; Teng et al., 2024). However, such methods have critical drawbacks. First, structure-based attacks inherently constrain the attack surface, resulting in predictable patterns that are easily mitigated by modern defense mechanisms (Wang et al., 2024a) and detectable by OCR systems. Second, such images are primarily designed to bypass sensitive text detection, rather than actively inducing harmful continuations, limiting their applicability to implicit toxicity tasks.
More recently, automated red-teaming frameworks have emerged to systematically uncover vulnerabilities in VLMs. For instance, IDEATOR (Wang et al., 2024c) introduced the ”VLMs against VLMs” paradigm, employing a red-team VLM to generate jailbreak prompts paired with images from general-purpose diffusion models (Rombach et al., 2022). While innovative, this approach inherits a fundamental limitation: the inherent conflict between diffusion models’ original training objective (faithful visual-textual alignment) and red teaming goals (toxicity maximization). Off-the-shelf diffusers like Stable Diffusion (Rombach et al., 2022) are optimized to avoid generating harmful content during pretraining, creating an intrinsic bias against jailbreaking effectiveness.
This highlights a critical yet unexplored research opportunity: Can we reconfigure diffusion models to actively induce toxicity while maintaining semantic plausibility? Current approaches treat image generation and toxicity induction as decoupled processes - first generating ”safe” images with pretrained diffusers, then hoping they coincidentally trigger harmful outputs. Our work overturns this passive paradigm through reinforced diffusion alignment, where the diffusion model is actively tuned to: (1) maximize toxicity through RL-driven reward signals; (2) preserve semantic coherence with image prompts; (3) dynamically adapt to target VLM’s defense patterns. By transforming diffusion models from general-purpose generators into active toxicity amplifiers, we establish a new paradigm for multimodal red teaming that fundamentally rethinks how visual modalities can be weaponized against VLMs.
In this work, we propose Red Team Diffuser (RTD), a novel framework that redefines multimodal jailbreak through semantically adversarial image synthesis. Departing from conventional static attack patterns, RTD introduces a dynamic two-phase optimization paradigm. First, we leverage the contextual reasoning capabilities of an LLM to perform a greedy search over image prompts, identifying those that maximize the toxicity of VLM responses and are thus highly effective for red teaming. Second, the discovered prompts drive RL-based fine-tuning of a diffusion model, using a dual-objective reward: a toxicity reward to amplify the harmfulness of the induced outputs, and an image-prompt alignment reward to ensure semantic consistency as a regularization term. This transforms standard diffusion models into toxicity-aware generators, fundamentally altering their output distribution toward vulnerable regions of target VLMs’ multimodal alignment space.
The main contributions of our work are threefold:
  • New Attack Paradigm: We propose Red Team Diffuser (RTD), the first framework to integrate reinforcement learning with diffusion models for semantically adversarial image generation, establishing a new paradigm for multimodal red teaming. Unlike prior approaches that rely on static patterns or unmodified diffusers, RTD actively co-optimizes both image prompts and diffusers for toxicity induction.
  • Technical Innovations: We employ an LLM to automatically generate diverse image prompts, followed by iterative toxicity evaluation to identify the most effective triggers. Furthermore, we introduce an RL-based fine-tuning strategy with dual rewards, balancing toxicity maximization and semantic preservation, thereby ensuring both high attack potency and natural image quality.
  • Empirical Contributions: We demonstrate a 10.69% increase in toxicity rate over text-only attacks on LLaVA. Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, along with toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s LLaMA. Our findings reveal a critical vulnerability in cross-modal toxicity amplification, where semantically adversarial images exacerbate text-based harm.
2. Related Work
2.1. Large Vision-Language Models
@@
Large Vision-Language Models (VLMs) extend Large Language Models (LLMs) by integrating visual modalities through an image encoder and a cross-modal alignment module. This design enables VLMs to perform instruction-following and open-ended reasoning over both text and images. Recent open-source VLMs have explored various architecture designs. LLaVA (Liu et al., 2024a) connects a CLIP encoder (Radford et al., 2021) with LLaMA (Touvron et al., 2023), using GPT-4 (Achiam et al., 2023) to synthesize multimodal instruction-following data for end-to-end fine-tuning. MiniGPT-4 (Zhu et al., 2023) employs a linear projection to align a frozen ViT (Dosovitskiy et al., 2020) with Vicuna (Chiang et al., 2023), using pretraining on large-scale image-text pairs followed by high-quality fine-tuning. InstructBLIP (Dai et al., 2023), based on BLIP-2 (Li et al., 2023b), introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding. While these models achieve impressive performance, the incorporation of visual inputs significantly broadens the attack surface compared to text-only LLMs. Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior (Shayegani et al., 2023b; Liu et al., 2024b; Li et al., 2023a; Ma et al., 2025). These findings highlight the urgent need to evaluate VLM robustness under multimodal threats—an area this work directly addresses.
Figure 2. Overview of the proposed method. Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation. The diffusion model generates an image based on this prompt. Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
2.2. Jailbreak Attacks against VLMs
@@
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks (Bagdasaryan et al., 2023; Bailey et al., 2023; Shayegani et al., 2023a; Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024; Wang et al., 2024b) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, making them less practical and more easily detectable (Nie et al., 2022; Zhang et al., 2023). In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR. For example, FigStep (Gong et al., 2023) converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP (Liu et al., 2023; Ma et al., 2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems. To address these limitations, the concept of red team VLMs has progressed with IDEATOR (Wang et al., 2024c), which autonomously generates jailbreak image-text pairs in a black-box setting. By offering a scalable and automated pipeline for discovering jailbreak prompts, IDEATOR demonstrates the promise of using VLMs themselves to expose alignment weaknesses. While IDEATOR focuses on prompt-level red teaming, our work introduces a fundamentally different and complementary attack paradigm: reinforced diffusion-based image optimization. Rather than relying on VLM-generated inputs, we investigate how semantically adversarial images—crafted via reinforcement learning—can effectively induce toxic text continuations when paired with fixed text prefixes. This setup presents a significantly more challenging threat model: it removes reliance on typographic or prompt engineering tricks and explores the vulnerability of VLMs to unrestricted, natural-looking visual stimuli. By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks, highlighting a critical blind spot in existing alignment strategies.
3. Preliminary
Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm (Black et al., 2023) to fine-tune a pretrained diffusion model (Rombach et al., 2022). Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP). At each step, actions (denoising predictions) are chosen to maximize the cumulative reward. The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps. Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$. The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process. The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$. The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$. That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
@@ 3. Preliminary
$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right].$
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator (Kakade and Langford, 2002):
$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg{[}\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg{]},$
@@
4. Methodology
4.1. Threat Model
Attack Goals  We focus on the sentence continuation task, an underexplored area in alignment mechanisms that poses unique challenges due to the open-ended nature of model outputs. The attacker aims to bypass the model’s safety measures and induce toxic continuations, such as the generation of obscene or profane language. Toxic continuations can also be viewed as a form of multimodal prompt injection, where carefully designed text prefixes and image triggers manipulate the model’s output into generating harmful content.
Adversary Capabilities  We assume the attacker has only black-box access to the victim VLM, simulating real-world interactions via external interfaces. By reinforcement fine-tuning red team diffusers based on the VLM’s feedback, RTD learns the VLM’s behavior patterns and identifies vulnerabilities, refining its policy to trigger toxic content.
4.2. Greedy Prompt Search
The effectiveness of our toxic continuation attacks hinges on identifying image prompts that maximally amplify toxicity in VLM-generated outputs. To overcome the limitations of directly using LLM-generated prompts, we introduce a dynamic, feedback-driven search procedure designed to discover high-risk image prompts tailored to specific toxic text prefixes. We employ Google’s Gemini Pro (Team et al., 2023)—with safety settings disabled—as a red team LLM to propose candidate image prompts conditioned on each text prefix from the LLM safety benchmark. Each image prompt is required to retain the semantic context of the prefix while subtly amplifying its harmful intent, thereby increasing the likelihood of inducing toxic completions. To enhance control over prompt generation, we apply in-context learning (Brown, 2020) by providing exemplars of previously identified harmful sentence continuations in the system prompt, which serve as implicit guidance for Gemini to produce image prompts consistent with known harmful patterns.

### v4

# RedDiffuser: Red Teaming Vision-Language Models for Toxic Continuation via Reinforced Stable Diffusion

@@
Abstract
Large Vision-Language Models (VLMs) remain vulnerable to jailbreak attacks that bypass safety mechanisms and elicit harmful outputs. In this work, we examine a subtle yet critical variant of this threat: toxic continuation. Unlike standard jailbreaks driven by malicious instructions, toxic continuation arises when VLMs are exposed to harmful prefixes paired with adversarial images, leading to toxic completions. To address this, we propose RedDiffuser (RedDiff), the first multimodal red teaming framework that leverages reinforcement learning to fine-tune diffusion models for generating natural-looking adversarial images. RedDiffuser combines greedy search to identify image prompts that maximize toxicity with reinforcement optimization that balances two objectives: (1) promoting toxic continuation, guided by a toxicity classifier, and (2) maintaining semantic coherence and stealthiness, measured by text–image similarity scores. This design circumvents traditional noise-based adversarial patterns and yields more realistic attack vectors. Experiments show that RedDiffuser significantly increases the toxicity rate of LLaVA-1.5 outputs by 10.69% and 8.91% on the original and hold-out sets, respectively. Moreover, it demonstrates strong transferability, raising toxicity rates by 5.1% on Gemini-1.5-flash and 26.83% on Llama-3.2-11B-Vision-Instruct. We also consider a more realistic deployment setting involving external NSFW guardrails. Incorporating a checker-aware training signal allows RedDiffuser to retain attack effectiveness while achieving high guardrail pass rates, indicating that toxic continuation vulnerabilities persist under safety-filtered conditions. These findings uncover two overlooked flaws in current VLM alignment: the failure to mitigate toxic continuation and the vulnerability to cross-modal adversarial amplification. Our results highlight the urgency of multimodal red teaming in evaluating VLM safety. We will release the RedDiffuser codebase to support future research. Disclaimer: This paper contains content that may be disturbing or offensive.
I Introduction
The rapid deployment of Vision-Language Models (VLMs) such as GPT-4o [achiam2023gpt] and Gemini [team2023gemini] has highlighted significant vulnerabilities in their safety alignment. Recent research demonstrates that VLMs are susceptible to jailbreak attacks, where adversaries can bypass built-in safety mechanisms and elicit harmful outputs. Among these, a more insidious threat emerges: toxic continuation. Despite progress in refusing explicit harmful instructions, existing defenses fall short in addressing this specific vulnerability, where the model, when presented with both a malicious input and an initial segment of harmful output, can be manipulated to generate highly toxic completions. This vulnerability is particularly pronounced in the multimodal setting, where images can substantially amplify the risk of toxic continuation (Figure 1). These gaps highlight the urgent need for systematic investigation and mitigation of this emerging risk, moving beyond the conventional focus on single-modality and instruction-based jailbreaks.
However, a critical challenge in addressing this multimodal vulnerability is identifying the very images that induce such toxic responses, a problem that remains largely unsolved. Li et al. [li2024images] show that even a blank image can sharply increase the success rate of jailbreak attacks, demonstrating that the connection between image content and VLM toxicity can defy intuitive understanding. While some white-box approaches employ gradient-based optimization to generate adversarial images [qi2024visual, wang2024white], these methods typically require full access to the model and often result in images with visible artifacts that are unnatural and easily detected [nie2022diffusion, zhang2023mutation]. Furthermore, their limited transferability across different VLMs [schaefferfailures] severely restricts their real-world applicability.
Figure 1: Example of toxic continuation on LLaVA. RedDiffuser generates images capable of inducing harmful continuations, whereas images from a standard diffusion model lead to benign VLM output.
These challenges highlight the need for stealthy, black-box attacks capable of generating natural-looking yet adversarial images without requiring access to model internals. Most existing black-box jailbreaks depend on hand-crafted pipelines or unmodified diffusion models [gong2023figstep, liu2023query], which constrain the range of possible attacks and focus mainly on evading text-based filters rather than proactively inducing toxic model responses. Recent methods such as IDEATOR [wang2024ideator] present a “VLMs against VLMs” approach, pairing prompts with images generated by off-the-shelf diffusion models. However, these diffusion models are typically pretrained to improve aesthetic quality [rombach2022high], making them inherently misaligned with the objectives of red teaming. This gap leads to a fundamental question: Can we directly optimize diffusion models to actively induce toxic continuations in VLMs while preserving image naturalness?
We address this challenge by introducing RedDiffuser (RedDiff), the first framework to employ reinforcement learning for red teaming with diffusion models. RedDiffuser operates in two coordinated phases. In the first phase, an LLM conducts a greedy search over candidate image prompts to identify those that maximize toxicity in the VLM’s output. These selected prompts then serve as the basis for the second phase, where the diffusion model is reinforcement fine-tuned with a dual reward: one term encourages toxic output (measured by Detoxify), while the other promotes semantic coherence (measured by BERTScore). This process turns the diffusion model from a neutral generator into an adversarial tool for exposing alignment vulnerabilities in VLMs. We also extend RedDiffuser to handle scenarios where external NSFW guardrails are present, introducing a checker-aware variant. Our main contributions are as follows:
  • We present RedDiffuser, the first framework to integrate reinforcement learning with diffusion models for generating semantically adversarial images, thereby enabling effective cross-modal jailbreaks.
  • Our approach combines greedy prompt search with reinforcement fine-tuning, guided by both toxicity and alignment rewards, to ensure strong attack efficacy while maintaining natural image quality.
  • RedDiffuser increases toxicity rates on LLaVA by 10.69% over text-only attacks, achieves 8.91% on a hold-out set, and transfers successfully to Gemini (+5.1%) and LLaMA-Vision (+26.83%).
  • Under external NSFW guardrails, the checker-aware variant of RedDiffuser improves the guardrail pass rate by +40.4% over Stable Diffusion, indicating that the attack remains feasible despite safety filtering.
II Related Work
II-A Vision-Language Models
@@
Large Vision-Language Models (VLMs) extend Large Language Models (LLMs) by integrating visual modalities through an image encoder and a cross-modal alignment module. This design enables VLMs to perform instruction-following and open-ended reasoning over both text and images. Recent open-source VLMs have explored various architecture designs. LLaVA [liu2024visual] connects a CLIP encoder [radford2021learning] with LLaMA [touvron2023llama], using GPT-4 [achiam2023gpt] to synthesize multimodal instruction-following data for end-to-end fine-tuning. MiniGPT-4 [zhu2023minigpt] employs a linear projection to align a frozen ViT [dosovitskiy2020image] with Vicuna [chiang2023vicuna], using pretraining on large-scale image-text pairs followed by high-quality fine-tuning. InstructBLIP [dai2023instructblip], based on BLIP-2 [li2023blip], introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding. While these models achieve impressive performance, the incorporation of visual inputs significantly broadens the attack surface compared to text-only LLMs. Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior [shayegani2023survey, liu2024survey, li2023privacy, ma2025safety]. These findings highlight the urgent need to evaluate VLM robustness under multimodal threats—an area this work directly addresses.
II-B Jailbreak Attacks on VLMs
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks [bagdasaryan2023ab, bailey2023image, shayegani2023jailbreak, carlini2024aligned, qi2024visual, niu2024jailbreaking, wang2024white] typically leverage gradient information to perturb input images or text, thereby increasing the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, which makes them less practical and easier to detect [nie2022diffusion, zhang2023mutation].
In contrast, black-box jailbreak attacks provide more practical and scalable solutions because they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite these advantages, many existing black-box approaches rely heavily on typography-based manipulations, which are vulnerable to detection techniques such as OCR. For example, FigStep [gong2023figstep] converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP [liu2023query, ma2024visual] also incorporate typography-driven attacks, reducing their stealthiness against advanced detection systems.
@@ II-B Jailbreak Attacks on VLMs
To overcome these limitations, the concept of red team VLMs has advanced with IDEATOR [wang2024ideator], which autonomously generates jailbreak image-text pairs in a black-box setting. By offering a scalable and automated pipeline for discovering jailbreak prompts, IDEATOR demonstrates the promise of using VLMs themselves to expose alignment weaknesses. While IDEATOR focuses on prompt-level red teaming, our work introduces a fundamentally different and complementary attack paradigm: reinforced diffusion-based image optimization. Instead of relying on VLM-generated inputs, we investigate how semantically adversarial images crafted via reinforcement learning can induce toxic text continuations when paired with fixed text prefixes. This setup constitutes a significantly more challenging threat model, as it removes dependence on typographic or prompt engineering tricks and examines the vulnerability of VLMs to unrestricted, natural-looking visual stimuli. By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks and highlight a critical blind spot in existing alignment strategies.
Figure 2: RedDiffuser overview. Given an incomplete toxic sentence, Gemini selects an image prompt via greedy search. A diffusion model generates an image, which is passed to a VLM (LLaVA) to produce a continuation. Toxicity and alignment scores from Detoxify and BERTScore are used as rewards to fine-tune the diffusion model.
III Preliminary
[…]
@@ III Preliminary
$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right]$ (1)
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator [kakade2002approximately]:
$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg[\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg]$ (2)
@@
IV Methodology
IV-A Threat Model
Attack Goals.  We target the sentence continuation task, where the model generates open-ended completions from partial inputs. The attacker’s objective is to bypass safety mechanisms and induce toxic content, including profanity, insults, or threats, by pairing adversarial images with fixed text prefixes. This constitutes a form of multimodal prompt injection, where carefully crafted image-text inputs bias the model toward unsafe responses. Adversary Capabilities.  We consider a black-box setting where the attacker interacts with the VLM solely through input-output queries, as in real-world API access. By fine-tuning the diffusion model via reinforcement learning with feedback from the VLM, the adversary learns to exploit alignment weaknesses and increase the likelihood of toxic completions.
IV-B Greedy Prompt Search
Toxic continuation attacks require identifying image prompts that can consistently amplify toxicity in VLM outputs. However, directly using prompts generated by a general LLM is often suboptimal. To overcome this limitation, we design a feedback‑driven greedy prompt search procedure that iteratively refines image prompts to maximize toxicity amplification for given toxic text prefixes. In our approach, we employ Google’s Gemini Pro [team2023gemini] with safety settings disabled as the red‑team LLM to propose candidate image prompts. To guide Gemini’s output toward semantically consistent yet harmful prompts, we utilize in-context learning [brown2020language] by providing Gemini with exemplars of previously identified harmful continuations.